the class average on a test is 85, with a standard deviation of 3.8, find the probability that a student received at least 76 on the test
Answer by stanbon(75887)

the class average on a test is 85, with a standard deviation of 3.8, find the probability that a student received at least 76 on the test
------------------
z(76) = (76-85)/3.8 = -2.3684
---
P(x>= 76) = P(z>= -2.3684) = 0.9911
=========================================
Cheers,
Stan H.